# Voltage dependent first-principles barriers to Li transport within Li-ion battery Solid-Electrolyte-Interphases

### v1

@@ Abstract
Charging a Li-ion battery requires Li ion transport between the cathode and the anode. This Li-ion transport is dependent upon (among other factors) the electrostatic environment the ion encounters within the Solid-Electrolyte-Interphase (SEI), which separates the anode from the surrounding electrolyte. Previous first principles work has illuminated the reaction barriers through likely atomistic SEI environments, but has had difficulty accurately reflecting the larger electrostatic potential landscape that an ion encounters moving through the SEI. In this work, we apply the recently developed Quantum Continuum Approximation (QCA) technique to provide an equilibrium electronic potentiostat for first-principles interface calculations. Using QCA, we calculate the potential barrier for Li-ion transport through LiF, LiO, and LiCO SEIs along with LiF-LiF, and LiF-LiO grain boundaries, all paired with Li metal anodes. We demonstrate that the SEI potential barrier is dependent on the anode electrochemical potentials in each system. Furthermore, we find that the threshold potential at which the SEI potential barrier switches from encouraging Li ion transport toward the anode versus discouraging transport toward the anode is highly dependent on the exact SEI chemistry. Finally, we use these techniques to estimate the change in the diffusion barrier for a Li ion moving in a LiF SEI as a function of anode potential.
@@ 1 Introduction
Li-ion battery performance is crucial to meeting the world’s growing needs to decarbonize transportation 1. In addition to low weight and volume requirements, faster charging is an important factor in rendering electric vehicles competitive across broader use cases 2. The rate of charging is determined by a number of factors in the overall design of the battery cell, but is fundamentally determined by the speed at which Li-ions can be transported from the cathode to the anode 3. The transport of an individual Li ion is dependent on both the atomistic and electrostatic environment it sees at any given location. These can vary significantly across the cell, with one notable bottleneck being Li-ion transport through the (typically insulating) Solid Electrolyte Interphase (SEI) layer that develops between the anode and the electrolyte 4. It has been experimentally difficult to characterize the SEI at sufficient detail to understand the precise atomic environment experienced by Li ions.
First-principles computational work has been crucial to developing an understanding of Li-ion transport through SEIs 5, 6. Due to computational limits, the majority of work in this field has focused not directly on the interface, but on finding reaction barriers in bulk versions of common SEI’s such as LiF, LiO, and LiCO 7. More recent work has begun to focus on the direct interface between the anode and the SEI, finding that this interface can be the site of significant changes to the reaction barriers compared to bulk and is crucial for a full understanding of battery behavior 8, 9. Additional work has also examined the role of grain boundaries in the SEI, finding that these play a potentially significant role in reducing Li-ion transport barriers and thus likely serving as the main route for Li transfer 10, 11, 12.
While this work has significantly addressed the impact of the atomistic environment on Li-ion transport, the electrostatic potential environment has been less thoroughly examined. As demonstrated in Fig. 1, the electrostatic potential experienced by a given Li ion is dependent on its location within the extended battery interface. This potential difference across the length of the SEI can give rise to a significant potential barrier that a charge carrier must travel through to move from the electrolyte to the anode. The exact height and spatial location of the barrier are significantly modified by interfacial atomic interactions between the anode and SEI, which need quantum mechanical determination for each specific interface and voltage condition 13, making the height and distribution of barriers inherently voltage dependent. These interfacial barriers have indeed been observed to change with voltage in solid state battery experiments 14, 15, but remain largely unexplored in non-solid state batteries due to the difficulty of in-situ characterization.
@@
In this work, we use the recently developed Quantum Continuum Approximation (QCA) technique 16 to predict the first-principles potential barriers for Li ion transport as a function of both the atomistic and electrostatic potential environment. QCA works by coupling explicit DFT calculations of relevant interfaces (e.g. anode-SEI, SEI-electrolyte) to Poisson-Boltzmann distributions of charge in the bulk insulating SEI region. This allows us to extract potential barriers at electric fields relevant to battery charging operation and assign equilibrium electronic voltages to our DFT calculations. We examine the barrier for Li ion transport through LiF, LiO, and LiCO SEI environments as well as LiF-LiF, and LiF-LiO grain boundaries with a metal Li anode. We demonstrate that these interfacial barriers are highly dependent on the voltage at which the battery is being charged as well as the atomic interactions of the interface. We then examine the impact of these SEI barriers on the reaction barrier magnitude of a Li ion moving a single lattice site via the knock off mechanism in a LiF SEI as a function of voltage. Our results suggest altering the structure of the anode-SEI interface and controlling electronic voltage as useful tools for aiding the transport of Li ions in batteries and thus enhancing charging/discharging behavior.
@@ 2.1 Determining Potential Gradients across battery SEIs
where is the applied voltage. We can then finally determine the potential dependent SEI barrier as
Note that the interfacial potential is also potential dependent.
@@ 2.2 The Quantum Continuum Approximation (QCA) for determining voltage on a Li ion SEI interface
To calculate the first principles response of Li ion transport in Li-ion batteries, we use QCA for metal-insulator-electrolyte interfaces, the full details of which have been previously described by Campbell 16. This QCA methodology was first developed for semiconducting-electrolyte interfaces 22, and has demonstrated excellent agreement with experimental characterization for Schottky barriers in metal-semiconductor interfaces 23.
In brief, explicit DFT calculations of both the metal-insulator and insulator-electrolyte interfaces along with regions of interest, such as atoms surrounding Li ion transport, are coupled to continuum representations of the Poisson-Boltzmann charge-voltage distribution within a bulk insulator. To incorporate the impacts of voltage, a bulk total charge for the entire electrode is selected, and then multiple distributions of charge between the different interfaces and the bulk insulator are calculated. Each of these differing charges introduces different electric fields to each section. We calculate many different charge distributions to find the one that causes the electric field to be continuous across the entire interface (as in Fig. 1) and the Fermi level to be in equilibrium throughout each system (as in Fig. 2b). From this we can then use the Trassati relation 24 to extract an equilibrium electronic anode potential for the system as
where is the electronic charge, assuming that the potential far away from the electrode has been set to zero. This offset of 1.37 is based on an aqueous electrolyte and is somewhat approximate. represents the electrochemical potential of one of of the halves of our battery system, in the case of our examples, the anode. This anode potential can be related to the open circuit voltage of the battery system by measuring the difference between the cathode potential and the anode potential ,
We note that this is an electronic definition of voltage, which is typically used throughout the computational electrochemistry community 25. This differs from the widespread definition of ionic voltage within the battery DFT community 26, which is based on the energy of inserting or removing Li atoms from the simulation
@@ 2.3 QCA for voltage dependent diffusion barriers
To calculate how the electronic potential impacts the diffusion barrier of a given Li ionic transport reaction, we first calculate the transition state using standard DFT Nudged Elastic Band (NEB) methodology. We then calculate how the energy of the both the transition state and the starting state of the reaction change as a function of voltage using QCA. The electronic potential specific diffusion barrier at a specific electronic potential can thus be calculated as
where and are the DFT calculated energy of the transition state and initial state, respectively, using QCA to control for the electronic potential. This approach mimics a similar approach recently taken by Vijay et al. to calculate potential dependent reaction barriers for molecular catalysis on metallic surfaces 27. Due to computational restrictions, we assume that the transition state remains the same across different potentials, which is largely consistent with the results from Vijay et al.27
@@ 2.4 QCA and Computational Details
The exact thickness and defect concentration of a given SEI are highly dependent on the anode, the electrolyte environment, and cycling conditions 28. Throughout this work, we will assume a temperature of 300 K, a SEI thickness of 200 Å, and a charge carrier concentration of cm as representative values 7, 29. The exact voltages and electric fields reported in this work and their impact on Li-ion transport throughout the SEI are necessarily dependent on these values. However, since we are working with relatively low charge carrier concentrations (at least compared to typical semiconducting systems), we do not see significant band bending as might be expected in a more semiconducting system (at least on the length scale of a few hundred Å). Instead, the potential continues in an essentially straight line throughout the SEI, following the slope established at the surface until meeting the secondary interface with the metal. We thus see little change in the reported barriers from increasing or decreasing the charge carrier concentration by a few orders of magnitude. In this paper, we will solely examine cases where the anode is a pure Li metal, but our methodology is broadly applicable to any anode and we expect that the general conclusions about the dependence of Li-ion transport barriers on the electronic potential and SEI chemistry of the battery will remain largely the same.
@@
All electronic structure calculations are done using the quantum espresso package 30. We use norm-conserving pseudopotentials from the PseudoDojo repository 31 and the Perdew-Burke-Ernzerhof exchange-correlation functional 32. We use kinetic energy cutoffs of 50 Ry and 400 Ry for the plane wave basis sets used to describe the Kohn-Sham orbitals and charge density, respectively. We use a 221 Monkhorst-Pack grid 33 to sample the Brillioun zone in our calculations. Interface structrues were generated using the pymatgen interface generator 34. Grain boundary structures are generated using the AIMSGB algorithm and toolset 35. Equilibrium structures are found by allowing the forces to relax below 0.05 eV/Å.
@@
We use the Environ package 36 to calculate the parabolic corrections to the surface dipole as well as including the planes of Helmholtz charges needed for QCA. For simplicity, all calculations are done in vacuum with only a first monolayer of dioxolane (DOL) molecules included as part of the electrolyte, which has been shown to improve band alignment of DFT calculations 37. Exact atomic coordinates and structural parameters used for all calculations are included in the supporting information.
@@ 3.1 SEI barriers
Using the QCA methodology established in Sec. 2, we simulate the SEI potential barrier as a function of voltage in LiF, LiO, LiCO, as well as a LiF-LiF and LiF-LiO grain boundary SEI systems.
@@ 3.1.1 LiF
We first examine the SEI barrier of LiF, where the interfacial anode-SEI structure is shown in Fig. 3a and the SEI-electrolyte structure shown in Fig. 3b. We calculate both the SEI barrier that a Li ion will face moving from the electrolyte through the SEI to the anode surface as well as the interfacial barrier a Li ion will need to overcome in moving through the anode-SEI interface in Fig. 3c. Both these values are dependent on the overall voltage of the system and are zero at the ’flatband’ voltage of the system, which in this case corresponds to 1.88 V (Li/Li). Away from the flatband voltage, the barrier is linearly responsive to the change in voltage, as will be true for all of the systems examined in this work.
A linear relationship between the SEI barrier and the voltage is not guaranteed for any given system, and is dependent on the extent of the band bending within the SEI. In this case, we are assuming a charge carrier concentration of 10 cm and a length of 20 nm, which, while relatively typical in batteries, are both relatively low relative to the typical semiconductor interface. This means that the electrostatic potential ends up exhibiting essentially no band bending within the SEI, and the potential profile in space is, in fact, linear. This constant slope of the potential within the SEI results in a linear relationship between the voltage and the potential for these systems. A higher charge concentration (e.g. 10 cm) and/or a longer length SEI would result in a more pronounced band bending, which would partially break this linear relationship between the voltage and the SEI barrier magnitude.
@@
From the perspective of a Li ion moving from the electrolyte to the anode at a given electronic voltage, it will have to overcome (or be helped by) the SEI potential barrier over the 20 nm of the bulk SEI width. We take the convention that a positive SEI barrier implies that the Li ion will have to overcome the potential barrier, and a negative barrier implies that the Li ion will be able to move “downhill” with respect to the electrostatic potential, i.e., be encouraged to move toward the anode. Given the positive slope of the SEI barrier as a function of voltage, our results imply that it will become easier and easier for the Li ion to migrate to the anode as the anode electrochemical potential decreases. This matches experimental findings of Li plating becoming stable as the anode electrochemical potential decreases 38. It further supports naive intuitions of battery charging becoming easier at higher open circuit battery voltages , as can be seen by lowering the anode electrochemical potential in Eq. 6. Notably, the point at which the SEI barrier encourages Li migration versus discouraging it is highly interface specific (in the case of the Li-LiF structure studied here, it occurs at 1.88 V (Li/Li)). We next go on to see how the relationship between these barriers changes with a differing SEI composition.
3.1.2 LiO

### v2

# Voltage dependent first-principles barriers to Li transport within Li ion battery Solid Electrolyte Interphases

@@ Abstract
Charging a Li ion battery requires Li ion transport between the cathode and the anode. This Li ion transport is dependent upon (among other factors) the electrostatic environment the ion encounters within the Solid Electrolyte Interphase (SEI), which separates the anode from the surrounding electrolyte. Previous first principles work has illuminated the reaction barriers through likely atomistic SEI environments, but has had difficulty accurately reflecting the larger electrostatic potential landscape that an ion encounters moving through the SEI. In this work, we apply the recently developed Quantum Continuum Approximation (QCA) technique to provide an equilibrium electronic potentiostat for first-principles interface calculations. Using QCA, we calculate the potential barrier for Li ion transport through LiF, Li2O, and Li2CO3 SEIs along with LiF-LiF, and LiF-Li2O grain boundaries, all paired with Li metal anodes. We demonstrate that the SEI potential barrier is dependent on the anode electrochemical potentials in each system. Finally, we use these techniques to estimate the change in the diffusion barrier for a Li ion moving in a LiF SEI as a function of anode potential. We find that properly accounting for interface and electronic voltage effects significantly lowers reaction barriers compared to previous literature results.
@@ 1 Introduction
Li ion battery performance is crucial to meeting the world’s growing needs to decarbonize transportation 1. In addition to low weight and volume requirements, faster charging is an important factor in rendering electric vehicles competitive across broader use cases 2. The rate of charging is determined by a number of factors in the overall design of the battery cell, but is fundamentally determined by the speed at which Li ions can be transported from the cathode to the anode 3. The transport of an individual Li ion is dependent on both the atomistic and electrostatic environment it sees at any given location. These can vary significantly across the cell, with one notable bottleneck being Li ion transport through the (typically insulating) Solid Electrolyte Interphase (SEI) layer that develops between the anode and the electrolyte 4. It has been experimentally difficult to characterize the SEI at sufficient detail to understand the precise atomic environment experienced by Li ions.
First-principles computational work has been crucial to developing an understanding of Li ion transport through SEIs 5, 6. Due to computational limits, the majority of work in this field has focused not directly on the interface, but on finding reaction barriers in bulk versions of common SEI’s such as LiF, Li2O, and Li2CO3 7. More recent work has begun to focus on the direct interface between the anode and the SEI, finding that this interface can be the site of significant changes to the reaction barriers compared to bulk and is crucial for a full understanding of battery behavior 8, 9. Additional work has also examined the role of grain boundaries in the SEI, finding that these play a potentially significant role in reducing Li ion transport barriers and thus likely serving as the main route for Li transfer 10, 11, 12.
While this work has significantly addressed the impact of the atomistic environment on Li ion transport, the electrostatic potential environment has been less thoroughly examined. As demonstrated in Fig. 1, the electrostatic potential experienced by a given Li ion is dependent on its location within the extended battery interface. This potential difference across the length of the SEI can give rise to a significant potential barrier that a charge carrier must travel through to move from the electrolyte to the anode. The exact height and spatial location of the barrier are significantly modified by interfacial atomic interactions between the anode and SEI, which need quantum mechanical determination for each specific interface and voltage condition 13, making the height and distribution of barriers inherently voltage dependent. These interfacial barriers have indeed been observed to change with voltage in solid state battery experiments 14, 15, but remain largely unexplored in non-solid state batteries due to the difficulty of in-situ characterization.
@@
In this work, we use the recently developed Quantum Continuum Approximation (QCA) technique 16 to predict the first-principles potential barriers for Li ion transport as a function of both the atomistic and electrostatic potential environment. QCA works by coupling explicit DFT calculations of relevant interfaces (e.g. anode-SEI, SEI-electrolyte) to Poisson-Boltzmann distributions of charge in the bulk insulating SEI region. This allows us to extract potential barriers at electric fields relevant to battery charging operation and assign equilibrium electronic voltages to our DFT calculations. We examine the barrier for Li ion transport through LiF, Li2O, and Li2CO3 SEI environments as well as LiF-LiF, and LiF-Li2O grain boundaries with a metal Li anode. We demonstrate that these interfacial barriers are highly dependent on the voltage at which the battery is being charged as well as the atomic interactions of the interface. We then examine the impact of these SEI barriers on the reaction barrier magnitude of a Li ion moving a single lattice site via the knock off mechanism in a LiF SEI as a function of voltage. Our results suggest that properly accounting for voltage and interface effects may significantly lower calculated Li diffusion barriers.
@@ 2.1 Determining Potential Gradients across battery SEIs
where is the applied voltage. We can then finally determine the voltage dependent SEI barrier as
Note that the interfacial potential is also voltage dependent.
@@ 2.2 The Quantum Continuum Approximation (QCA) for determining voltage on a Li ion SEI interface
To calculate the first principles response of Li ion transport in Li ion batteries, we use QCA for metal-insulator-electrolyte interfaces, the full details of which have been previously described by Campbell 16. This QCA methodology was first developed for semiconducting-electrolyte interfaces 22, and has demonstrated excellent agreement with experimental characterization for Schottky barriers in metal-semiconductor interfaces 23.
In brief, explicit DFT calculations of both the metal-insulator and insulator-electrolyte interfaces along with regions of interest, such as atoms surrounding Li ion transport, are coupled to continuum representations of the Poisson-Boltzmann charge-voltage distribution within a bulk insulator. For a perfectly insulating/semiconducting system, the charge distribution can be well approximated by the following Poisson–Boltzmann relationship:24
To incorporate the impacts of voltage, a total charge for the entire electrode is selected, and then multiple distributions of charge between the anode, SEI, and surface are calculated. To balance these included charges, we use Helmholtz planes of countercharge. Each of these differing charges introduces different electric fields to each section distinct DFT interface calculation. We can then quantify the potential distribution of the bulk SEI region using the Poisson-Boltzmann conditions outlined in Eq. 5. The potential within the SEI region can then be approximated as:
where is the planar averaged potential in the system due to the inclusion of voltage/charge, is the Boltzmann constant, and is the temperature of the system. This relationship applies throughout the bulk oxide region, which lasts from to . With Eq. 6, we can then predict the voltage offset within the SEI. We then find the equilibrium distribution of charge by finding the distribution of charge between the anode, SEI, and surface that causes the Fermi level to stay constant throughout the entire system:
where is the DFT Fermi level of the SEI-solution interface, and is the DFT Fermi level of the metal anode-SEI interface. Since the countercharges within the model take the form of Helmholtz planes, this gives a Helmholtz model for the electrical double layer within the electrolyte. This provides a useful first approximation of the electrical double layer response, but in the future, this could be replaced with a Guoy-Chapman or similar improved model.
From this we can then use the Trassati relation 25 to extract an equilibrium electronic anode potential for the system as
where is the electronic charge, assuming that the potential far away from the electrode has been set to zero. This offset of 1.37 is based on the standard electrode potential for a Li/Li+ reaction of a Li anode in an aqueous solution as compared to SHE26. In principle, the exact chemical makeup of the electrolyte would change this value, but we use this value as a useful first approximation, particularly for interpreting results relative to each other. represents the electrochemical potential of one of of the halves of our battery system, in the case of our examples, the anode. This anode potential can be related to the open circuit voltage of the battery system by measuring the difference between the cathode potential and the anode potential ,
We note that this is an electronic definition of voltage, which is typically used throughout the computational electrochemistry community 27. This differs from the widespread definition of ionic voltage within the battery DFT community 28, which is based on the energy of inserting or removing Li atoms from the simulation
@@ 2.3 QCA for voltage dependent diffusion barriers
To calculate how the electronic potential impacts the diffusion barrier of a given Li ionic transport reaction, we first calculate the transition state using standard DFT Nudged Elastic Band (NEB) methodology. This methodology generated several images along the transition pathway from the initial to the final state. This transition pathway is then optimized and the image with the highest energy, termed the climbing image, is found and the reaction barrier calculated as the difference between the climbing state energy and the initial state energy. With the climbing image identified, we calculate how the energy of the both the climbing image and the starting state of the reaction change as a function of voltage using QCA. The electronic potential specific diffusion barrier at a specific electronic potential can thus be calculated as
where and are the DFT calculated energy of the climbing image and initial state, respectively, using QCA to control for the electronic potential. This approach mimics a similar approach recently taken by Vijay et al. to calculate potential dependent reaction barriers for molecular catalysis on metallic surfaces 29. Due to computational restrictions, we assume that the climbing image remains the same across different potentials, which is largely consistent with the results from Vijay et al.29
@@ 2.4 QCA and Computational Details
The exact thickness and defect concentration of a given SEI are highly dependent on the anode, the electrolyte environment, and cycling conditions 30. Throughout this work, we will assume a temperature of 300 K, a SEI thickness of 200 Å, and a charge carrier concentration of cm-3 as representative values 7, 31. The exact voltages and electric fields reported in this work and their impact on Li ion transport throughout the SEI are necessarily dependent on these values. However, since we are working with relatively low charge carrier concentrations (at least compared to typical semiconducting systems), we do not see significant band bending as might be expected in a more semiconducting system (at least on the length scale of a few hundred Å). Instead, the potential continues in an essentially straight line throughout the SEI, following the slope established at the surface until meeting the secondary interface with the metal. We thus see little change in the reported barriers from increasing or decreasing the charge carrier concentration by a few orders of magnitude. These results are only applicable for the charge carrier concentration and SEI lengths we examine here. Due to computational restrictions, we cannot examine all possibilities; this does mean, however, that these results may not be representative all operational conditions of a battery. This issue is particularly exacerbated by the fact that the charge carrier concentration may not be uniform throughout a realistic SEI. Future work will focus on developing more localized sections of charge density and ionic pathways to reflect this mix. This would result in changes to Eq. 6 and the need for more explicit calculations at regions of interest. It may be reasonable to treat current results as representing an average charge carrier concentration across a broad area, however. In this paper, we solely examine cases where the anode is a pure Li metal, but our methodology is broadly applicable to any anode and we expect that the general conclusions about the dependence of Li ion transport barriers on the electronic potential and SEI chemistry of the battery will remain largely the same.
@@
All electronic structure calculations are done using the quantum espresso package 32. We use norm-conserving pseudopotentials from the PseudoDojo repository 33 and the Perdew-Burke-Ernzerhof exchange-correlation functional 34. We use kinetic energy cutoffs of 50 Ry and 400 Ry for the plane wave basis sets used to describe the Kohn-Sham orbitals and charge density, respectively. We use a 221 Monkhorst-Pack grid 35 to sample the Brillioun zone in our calculations. Interface structrues were generated using the pymatgen interface generator 36. Grain boundary structures are generated using the AIMSGB algorithm and toolset 37. Equilibrium structures are found by allowing the forces to relax below 0.05 eV/Å.
@@
We use the Environ package 38 to calculate the parabolic corrections to the surface dipole as well as including the planes of Helmholtz charges needed for QCA. For simplicity, all calculations are done in vacuum with only a first monolayer of dioxolane (DOL) molecules included as part of the electrolyte, which has been shown to improve band alignment of DFT calculations 39. In the supporting information, we demonstrate the dependence of the calculated work function on the number of DOL layers included, finding reasonable variability with the number of included DOL layers, although our monolayer results are within 0.2 eV of a three layer structure, indicating this is at least useful as a first approximation of the electrolyte system. Exact atomic coordinates and structural parameters used for all calculations are included in the supporting information.
@@ 3.1 SEI barriers
Using the QCA methodology established in Sec. 2, we simulate the SEI potential barrier as a function of voltage in LiF, Li2O, Li2CO3, as well as a LiF-LiF and LiF-Li2O grain boundary SEI systems. For each of these systems, we report the interfacial barrier as well, which, as outlined above, is key to determining the total potential drop across the SEI as a function of voltage, and thus for impacting Li ion transport through the SEI.
@@ 3.1.1 LiF
We first examine the SEI barrier of LiF, where the interfacial anode-SEI structure is shown in Fig. 3a and the SEI-electrolyte structure shown in Fig. 3b. We calculate both the SEI barrier that a Li ion will face moving from the electrolyte through the SEI to the anode surface as well as the interfacial barrier a Li ion will need to overcome in moving through the anode-SEI interface in Fig. 3c. Both these values are dependent on the overall voltage of the system and are zero at the ’flatband’ voltage of the system, which in this case corresponds to 1.88 V (Li/Li+). Away from the flatband voltage, the barrier is linearly responsive to the change in voltage, as will be true for all of the systems examined in this work.
A linear relationship between the SEI barrier and the voltage is not guaranteed for any given system, and is dependent on the extent of the band bending within the SEI. In this case, we are assuming a charge carrier concentration of 109 cm-3 and a length of 20 nm, which, while relatively typical in batteries, are both relatively low relative to the typical semiconductor interface. This means that the electrostatic potential ends up exhibiting essentially no band bending within the SEI, and the potential profile in space is, in fact, linear. This constant slope of the potential within the SEI results in a linear relationship between the voltage and the potential for these systems. A higher charge concentration (e.g. 1014 cm-3) and/or a longer length SEI would result in a more pronounced band bending, which would partially break this linear relationship between the voltage and the SEI barrier magnitude.
@@
From the perspective of a Li ion moving from the electrolyte to the anode at a given electronic voltage, it will have to overcome (or be helped by) the SEI potential barrier over the 20 nm of the bulk SEI width. We take the convention that a positive SEI barrier implies that the Li ion will have to overcome the potential barrier, and a negative barrier implies that the Li ion will be able to move “downhill” with respect to the electrostatic potential, i.e., be encouraged to move toward the anode. Given the positive slope of the SEI barrier as a function of voltage, our results imply that it will become easier and easier for the Li ion to migrate to the anode as the anode electrochemical potential decreases. This matches experimental findings of Li plating becoming stable as the anode electrochemical potential decreases 40. It further supports naive intuitions of battery charging becoming easier at higher open circuit battery voltages , as can be seen by lowering the anode electrochemical potential in Eq. 9. Notably, the point at which the SEI barrier encourages Li migration versus discouraging it is highly interface specific (in the case of the Li-LiF structure studied here, it occurs at 1.88 V (Li/Li+)). We next go on to see how the relationship between these barriers changes with a differing SEI composition.
3.1.2 Li2O